Consider the linear systemProblem 1. (10 points) Consider the linear system 4 = [-12 -8 a. Find the eigenvalues and eigenvectors for the coefficient

y⃗ ′=[6−124−8]y⃗ .
Answer #1

6-1 1-12 -8-1 * 9 F23] =45 Let t be the eigenvalue of 4. So, characterstic eqr. is: IA-d11=0 1 = (6-1) (-8-1) +48 = 0 12 +